# Galaxy Zoo: Cosmic Dawn – morphological classifications for over 41 000 galaxies in the Euclid Deep Field North from the Hawaii Two-0 Cosmic Dawn survey

### v1

@@ 6 Conclusion
The subjects in this iteration came from the Hawaii Twenty Square Degree survey as part of the wider Cosmic Dawn survey, featuring deep multiband Hyper Suprime-Cam (HSC) imaging from six square degrees of the Euclid Deep Field North, down to a depth of $m_{{\rm HSC}-i}=21.5$ for the purpose of resolving galaxy morphology. With an updated decision tree of questions, volunteers were asked to identify clumpy galaxies and tag rare objects like gravitational lenses and extremely red objects, which are of particular interest given the wide redshift range of this data set (up to $z_{\rm phot}=2.5$; median of 0.42). The final morphology catalogues in this data release contain a wide range of classified objects within the Euclid Deep Field North (EDFN), with suggested selection cuts providing a means of interrogating the results for a given science case, such as accessing a pure sample of edge-on disks, spirals, clumpy galaxies, and rarer irregular or ring galaxies. Given the sky coverage, this rich data set provides an opportunity to extract objects in the EDFN for further examination with deep Euclid imaging and high-resolution follow-up observations. For example, the data has already given rise to 51 newly discovered gravitational lenses after expert vetting of HSC and Euclid imaging of lens candidates found by volunteers. Additionally, with the ground-based HSC’s seeing comparable to that of the Vera C. Rubin Observatory, the classified images will act as a valuable truth set for training models to classify the wealth of data from the newly begun Legacy Survey of Space and Time.
@@
The Galaxy Zoo project allows for the visual inspection of galaxies by humans, in order to correctly classify their apparent morphologies and for the serendipitous discovery of new and interesting features that would be missed by machine learning algorithms. On the other hand, such work requires many (typically up to 40) volunteers per subject image for the classifications to be rigorous, and can therefore take a long time for a data set to be fully classified. Galaxy Zoo’s deep learning model, Zoobot, provides a two-stage method for circumventing this. First, Zoobot can separate out (retire) the easier subjects while leaving volunteers with more challenging and interesting subjects, whose classifications are then used to improve (fine-tune) Zoobot in an active learning cycle. Second, a fully fine-tuned Zoobot can confidently predict how volunteers would classify each image in a fraction of the time. Here, we find that the active learning cycle aided in retiring $\sim$78 per cent of subjects prior to them reaching 40 volunteer classifications, with the majority requiring less than 10 before Zoobot was confident enough to retire them. Then, applying a fully fine-tuned Zoobot to all subjects after project completion resulted in 68 per cent of vote fractions having mean deviations from volunteers below 0.2. Although larger than those in previous iterations, suggesting that further training is required on similar deep surveys, Zoobot’s predicted uncertainties were, on average, consistent with the discrepancies. As new iterations of Galaxy Zoo take place, the model will continue to improve and new volunteer classifications will help us explore galaxy evolution over ever more significant fractions of the age of the Universe.
@@ Acknowledgements
This publication has been made possible by the participation of more than 10 000 volunteers in the Galaxy Zoo project. GZCD is partly supported by the ESCAPE project, which aims to bring together the astronomy, astroparticle and particle physics communities to support open science, according to FAIR (Findable, Accessible, Interoperable and Reusable) principles. ESCAPE - The European Science Cluster of Astronomy & Particle Physics ESFRI Research Infrastructures has received funding from the European Union’s Horizon 2020 research and innovation programme under the Grant Agreement no. 824064. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan, and Princeton University.
JP and SS were partly supported by the ESCAPE project. JP, HD, and SS are supported by the ACME, ELSA, and OSCARS projects. ”ACME: Astrophysics Centre for Multimessenger studies in Europe”, ”ELSA: Euclid Legacy Science Advanced analysis tools”, and ”OSCARS: Open Science Clusters’ Action for Research and Society” are funded by the European Union under grant agreement no. 101131928, 101135203, and 101129751, respectively; ELSA is also funded by Innovate UK grant 10093177. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. LFF was partially supported by NSF award IIS-2334033. BDS acknowledges support through a UK Research and Innovation Future Leaders Fellowship [grant number MR/T044136/1] and its renewal [grant number MR/Z000076/1]. DBS gratefully acknowledges support from NSF Grant 2407752. IS acknowledges NASA grant 80NSSC24K1489. WJP has been supported by the Polish National Science Center project UMO-2023/51/D/ST9/00147.
This work made use of Astropy: [^5] http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration: Robitaille et al., 2013; Astropy Collaboration: Price-Whelan et al., 2018, 2022). Data analysis made use of the Python packages NumPy (Harris et al., 2020) and SciPy (Virtanen et al., 2020). Some figures were made with the Python package matplotlib (Hunter, 2007).
@@ Appendix A Example Classifications
Figure 20: Example GZCD subjects tagged by volunteers as having a “bright-red-companion”, which appear to contain extremely red point-source objects away from the cutout centre. Subjects are listed in order of their H20 ID, which is provided by each subject.
Figure 21: Example GZCD subjects tagged by volunteers as “beautiful”, “interesting”, “needs-more-research”, or similar tags. The majority correspond to spiral, irregular or disturbed galaxies with high angular resolution, as well as mergers, ring galaxies and clumpy galaxies. Subjects are listed in order of their H20 ID, which is provided by each subject.

### v2

@@ 6 Conclusion
The subjects in this iteration came from the Hawaii Twenty Square Degree survey as part of the wider Cosmic Dawn survey, featuring deep multiband Hyper Suprime-Cam (HSC) imaging from six square degrees of the Euclid Deep Field North, down to a depth of $m_{{\rm HSC}-i}=21.5$ for the purpose of resolving galaxy morphology. With an updated decision tree of questions, volunteers were asked to identify clumpy galaxies and tag rare objects like gravitational lenses and extremely red objects, which are of particular interest given the wide redshift range of this data set (up to $z_{\rm phot}=2.5$; median of 0.42). The final morphology catalogues in this data release contain a wide range of classified objects within the Euclid Deep Field North (EDFN), with suggested selection cuts providing a means of interrogating the results for a given science case, such as accessing a pure sample of edge-on disks, spirals, clumpy galaxies, and rarer irregular or ring galaxies. Given the sky coverage, this rich data set provides an opportunity to extract objects in the EDFN for further examination with deep Euclid imaging and high-resolution follow-up observations. For example, the data has already given rise to 51 newly discovered gravitational lenses after expert vetting of HSC and Euclid imaging of lens candidates found by volunteers. Additionally, with the ground-based HSC’s seeing comparable to that of the Vera C. Rubin Observatory, the classified images will act as a valuable truth set for training models to classify not only HSC imaging from the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS), but also the wealth of data from the newly begun Rubin Legacy Survey of Space and Time (LSST).
@@
The Galaxy Zoo project allows for the visual inspection of galaxies by humans, in order to correctly classify their apparent morphologies and for the serendipitous discovery of new and interesting features that would be missed by machine learning algorithms. On the other hand, such work requires many (typically up to 40) volunteers per subject image for the classifications to be rigorous, and can therefore take a long time for a data set to be fully classified. Galaxy Zoo’s deep learning model, Zoobot, provides a two-stage method for circumventing this. First, Zoobot can separate out (retire) the easier subjects while leaving volunteers with more challenging and interesting objects, whose classifications are then used to improve (fine-tune) Zoobot in an active learning cycle. Second, a fully fine-tuned Zoobot can confidently predict how volunteers would classify each image in a fraction of the time. Here, we find that the active learning cycle aided in retiring $\sim$78 per cent of subjects prior to them reaching 40 volunteer classifications, with the majority requiring less than 10 before Zoobot was confident enough to retire them. Then, applying a fully fine-tuned Zoobot to the whole subject set after project completion resulted in 68 per cent of vote fractions having mean deviations from volunteers below 0.2. Although larger than those in previous iterations, suggesting that further training is required on similar deep surveys, Zoobot’s predicted uncertainties were, on average, consistent with the discrepancies. As new iterations of Galaxy Zoo take place, the model will continue to improve and new volunteer classifications will help us explore galaxy evolution over ever more significant fractions of the age of the Universe.
@@ Acknowledgements
This publication has been made possible by the participation of more than 10 000 volunteers in the Galaxy Zoo project. We also gratefully acknowledge the contributions to this work by the Euclid Strong Lensing Working Group for their visual inspection efforts of the strong gravitational lens candidates, and we thank the anonymous referee for their useful comments that helped improve the paper.
GZCD is partly supported by the ESCAPE project, which aims to bring together the astronomy, astroparticle and particle physics communities to support open science, according to FAIR (Findable, Accessible, Interoperable and Reusable) principles. ESCAPE - The European Science Cluster of Astronomy & Particle Physics ESFRI Research Infrastructures has received funding from the European Union’s Horizon 2020 research and innovation programme under the Grant Agreement no. 824064. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan, and Princeton University.
JP and SS were partly supported by the ESCAPE project. JP, HD, and SS are supported by the ACME, ELSA, and OSCARS projects. ”ACME: Astrophysics Centre for Multimessenger studies in Europe”, ”ELSA: Euclid Legacy Science Advanced analysis tools”, and ”OSCARS: Open Science Clusters’ Action for Research and Society” are funded by the European Union under grant agreement no. 101131928, 101135203, and 101129751, respectively; ELSA is also funded by Innovate UK grant 10093177. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union. MW is a Dunlap Fellow. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. LFF was partially supported by the National Science Foundation (NSF) award IIS-2334033. BDS acknowledges support through a UK Research and Innovation Future Leaders Fellowship [grant number MR/T044136/1] and its renewal [grant number MR/Z000076/1]. DBS gratefully acknowledges support from NSF Grant 2407752. IS acknowledges NASA grant 80NSSC24K1489. WJP has been supported by the Polish National Science Center project UMO-2023/51/D/ST9/00147. This research was conducted under the agreement on joint mobility projects for the years 2024-2025 between the Polish Academy of Sciences and the National Science and Technology Council in Taiwan.
This work made use of Astropy: [^5] http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration, 2013, 2018, 2022). Data analysis made use of the Python packages NumPy (Harris et al., 2020) and SciPy (Virtanen et al., 2020). Some figures were made with the Python package matplotlib (Hunter, 2007).
@@ Appendix A Example Classifications
Figure 20: Example GZCD subjects tagged by volunteers as having a “bright-red-companion”, which appear to contain extremely red point-source objects away from the cutout centre. Subjects are listed in order of their H20 ID, which is provided next to each image.
Figure 21: Example GZCD subjects tagged by volunteers as “beautiful”, “interesting”, “needs-more-research”, or similar tags. The majority correspond to spiral, irregular or disturbed galaxies with high angular resolution, as well as mergers, ring galaxies and clumpy galaxies. Subjects are listed in order of their H20 ID, which is provided next to each image.